# Learning the Boundary of Solvability: Aligning LLMs to Detect Unsolvable Problems

### v1

@@ A.4.1 Game24 Generation
A.4.2 Hamiltonian Cycle and Path Generation
@@ A.4.3 Hitori Generation
A.4.4 Maze Navigation Generation
@@
To construct Unsolvable Instances, we employ a Strategic Blockage method rather than random wall placement. We analyze the solvable maze to identify all valid paths from S to E, and then select a Critical Junction—a coordinate $(x,y)$ that lies on a valid path (or is shared by multiple paths). We place an obstacle (wall) at $(x,y)$ and rigorously verify unsolvability using BFS. If the maze remains solvable (e.g., via a detour), we discard the instance or iteratively block remaining paths until connectivity is severed. This approach ensures that unsolvable mazes are structurally similar to solvable ones, often requiring the model to explore deep into the maze before realizing the path is blocked, thereby penalizing premature "give up" or hallucinated paths.
@@
A.5 Hyperparameters
We implement our training pipeline using the verl (HybridFlow) framework (sheng2025hybridflow), which supports efficient reinforcement learning. The specific hyperparameters used for Group Relative Policy Optimization (GRPO) are detailed in Table 4. We set the group size $G$ (denoted as rollout.n in verl) to 12. Notably, we explicitly set use_kl_loss=False, meaning the KL divergence penalty term was disabled during optimization, relying instead on the probability ratio clipping to maintain policy stability. We differentiate the batch size for Qwen3-1.7B and Qwen3-4B to accommodate their respective memory footprints.

### v2

@@ A.5.1 Game24 Generation
A.5.2 Hamiltonian Cycle and Path Generation
@@ A.5.3 Hitori Generation
A.5.4 Maze Navigation Generation
@@
To construct Unsolvable Instances, we employ a Strategic Blockage method rather than random wall placement. We analyze the solvable maze to identify all valid paths from S to E, and then select a Critical Junction, a coordinate $(x,y)$ that lies on a valid path (or is shared by multiple paths). We place an obstacle (wall) at $(x,y)$ and rigorously verify unsolvability using BFS. If the maze remains solvable (e.g., via a detour), we discard the instance or iteratively block remaining paths until connectivity is severed. This approach ensures that unsolvable mazes are structurally similar to solvable ones, often requiring the model to explore deep into the maze before realizing the path is blocked, thereby penalizing premature "give up" or hallucinated paths.
@@
A.6 Hyperparameters
We implement our training pipeline using the verl (HybridFlow) framework and vllm engine (sheng2025hybridflow; kwon2023efficient), which supports efficient reinforcement learning. The specific hyperparameters used for Group Relative Policy Optimization (GRPO) are detailed in Table 6. We set the group size $G$ (denoted as rollout.n in verl) to 12. Notably, we explicitly set use_kl_loss=False, meaning the KL divergence penalty term was disabled during optimization, relying instead on the probability ratio clipping to maintain policy stability.
We differentiate the batch size for Qwen3-1.7B and Qwen3-4B to accommodate their respective memory footprints. Specifically, the 1.7B model utilizes a larger batch size ($32\times 12$) compared to the 4B model ($14\times 12$). This adjustment balances the computational overhead, resulting in comparable training speeds for both architectures; typically, completing 120 training steps requires approximately one day on our 8 $\times$ NVIDIA A100 GPU cluster.